# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v2

@@ 2.5.2 Generative Feed-Forward View Synthesis
Regression-based methods work well for view interpolation, producing impressive visual results near the input views. However, it struggles with view extrapolation, leading to poor predictions for new views beyond the existing viewpoints, especially when estimating unseen regions of the scene. In contrast, generative feed-forward methods instead leverage generative models to synthesize realistic novel views based on learned data distributions, enabling view extrapolation even from a single input image. The generated multiview images can often be utilized as dense inputs for high-fidelity NeRF and 3DGS based reconstruction.
@@
Earlier works primarily use transformer-based autoregressive models [128, 152]. For example, GFVS [153] approaches novel view synthesis from a single view by sampling target images from a learned distribution conditioned on a source image and camera transformation, where the distribution is modeled autoregressively by leveraging a VQGAN [152] with a transformer. ViewFormer [154] further extends single-view NVS of GFVS to multiview NVS by VQVAE codebook [128] and transformer.
Recently, latent diffusion models [94] have been widely used in novel view synthesis due to their capability in generating high-resolution images, which encodes the input images into a latent space by a pretrained variational autoencoder and applies diffusion within the latent representation, and aims to learn the conditional distribution over the target image at the novel view. Diffusion-based generative methods have been widely explored using various diffusion priors, which will be introduced below.
@@
Image Diffusion Models. Zero-1-to-3 [121] leverages the latent diffusion model [94] pretrained for text-to-image generation and replaces text embedding with relative camera poses as conditioning to achieve novel view synthesis as illustrated in Fig. 7(a). ZeroNVS [155] extends Zero-1-to-3 to achieve single-view scene-level novel view synthesis by finetuning it on diverse large-scale object and real-scene datasets [27, 14, 156]. However, these methods still face challenges in generating consistent novel views. SyncDreamer [157] addresses this by initializing a diffusion model with Zero-1-to-3 weights and modeling the joint distribution of multiview images. Zero123++[12] arranges six views into a single image for improved multiview representation modeling. ViewDiff [158] also exploits diffusion priors, with emphasis on multi-view consistency in real-world data. Consistent123 [159] combines Zero-1-to-3 and stable diffusion to provide case-aware diffusion priors to ensure the multiview-consistency, while ConsistNet [160] performs parallel viewpoint-specific diffusion and aligns the generated images accordingly to enforce multiview geometric consistency. MVDream [89] proposes a multiview diffusion model that leverages both 2D and 3D data, combining the generalizability of 2D diffusion models with the consistency of 3D renderings. It further demonstrates that a multiview diffusion model can implicitly provide generalizable 3D priors without specific 3D representations.
@@
Video Diffusion Models. Video diffusion models [161, 162] have achieved impressive realism in video synthesis and are thought to inherently and implicitly capture 3D structures. Building on this capability, recent approaches have explored leveraging their priors to generate multiview images for high-quality 3D reconstruction. For example, ReconX [151], as illustrated in Fig. 7(b), harnesses the powerful generative prior of large pretrained video diffusion models [162] to synthesize novel views. It encodes extracted point clouds as 3D structural conditions, ensuring multiview consistency in the generated novel views. Similarly, ViewCrafter [8] first builds a point cloud representation and then uses point cloud renderings as the conditions of the video diffusion model [162] to enable consistent and accurate novel view synthesis. MultiDiff [163] leverages a single reference image and a predefined target camera trajectory as conditions, utilizing depth cues to encourage consistent novel view synthesis. More recently, GenFusion [164], DifFusion3D+ [165], and SpatialCrafter [166] have bridged reconstruction and generation by using video diffusion models as scene reconstruction refiners, enabling both artifact removal and scene content expansion.
@@ 3 Tasks & Applications
3.1 3D-aware Image Synthesis
3D-aware image synthesis refers to generating 2D images guided by an understanding of underlying 3D geometry, enabling the image synthesis of objects or scenes from different viewpoints while maintaining 3D consistency. These methods typically adopt a GAN-based framework, where a 3D representation is reconstructed in a feed-forward manner and then rendered to synthesize realistic 2D images.
Several earlier methods employ voxel-based representations (e.g., PlatonicGAN [167]) or 3D feature representations (e.g., HoloGAN [168] and BlockGAN [169]). However, these approaches often suffer from limited multiview consistency. To address this, GRAF [170] introduces a generative radiance field as a 3D representation, significantly improving consistency across different viewpoints. PiGAN [171] leverages implicit neural representations with periodic activation functions to model scenes as view-consistent radiance fields. Subsequently, GIRAFFE [172] constructs compositional generative radiance fields for scene representations, enabling controllable image synthesis. StyleNeRF [173] combines NeRF-based 3D representations with a style-based generative model for high-resolution, 3D-consistent image synthesis. EG3D [43] introduces an explicit-implicit triplane representation to achieve efficient and high-quality 3D-aware image synthesis.
Due to the high computational cost of volume rendering in implicit NeRF-based scene representations, Hyun et al. propose GSGAN [174], which replaces NeRF with 3D Gaussian Splatting (3DGS), enabling more efficient scene rendering through rasterization-based splatting. To stabilize the training of 3DGS-based 3D-aware image synthesis, GSGAN introduces hierarchical Gaussian representations, enabling coarse-to-fine scene modeling.
3.2 Camera-controlled Video Generation
To enable camera pose control in the video generation process, MotionCtrl [175], CameraCtrl [176], I2VControl-Camera [177] inject the camera parameters (extrinsic, Plücker embedding, or point trajectory) into a pretrained video diffusion model. Building upon this, CamCo [178] integrates epipolar constraints into attention layers, while CamTrol [179], NVS-Solver [180], and ViewExtrapolator [181] leverage explicit 3D point cloud renderings to guide the sampling process of the video diffusion models in a training-free manner. AC3D [182] carefully designs the camera representation injection to the pretrained model. ViewCrafter [8], Gen3C [183], and See3D [184] fine-tuned video diffusion models on point cloud renderings to enable better novel view synthesis. VD3D [185] enables camera control to transformer-based video diffusion models. Beyond static scenes, CameraCtrl II [186], and ReCamMaster [187] enable camera-controlled video generation on dynamic scenes by conditioning the video diffusion models on camera extrinsic parameters, while TrajectoryCrafter [188] also enables dynamic scene view synthesis by conditioning the video diffusion models on dynamic point cloud. Several recent works have advanced beyond single-camera scenarios: CVD [189], Vivid-ZOO [190], and SynCamMaster [191] develop frameworks for multi-camera synchronization.
3.3 Pose-free 3D Reconstruction
@@
The development of feed-forward models has enabled the reconstruction of 3D scenes from unposed images or videos without the need for per-scene optimization. FlowCam [192] uses a single-view feed-forward NeRF combined with optical flow to estimate poses and fuse multi-view point maps for NeRF-based 3D reconstruction. CoPoNeRF [193] builds and refines 4D correlation maps from image pairs to estimate flow and poses, enabling color and depth rendering.
@@
To extend these approaches into 3DGS, GGRt [95] employs PixelSplat [5] for predicting viewpoint-specific 3D Gaussian maps and introduces a pose estimation module that jointly optimizes camera poses alongside Gaussian predictions. PF3plat [103] proposes a coarse-to-fine strategy, estimating depth, confidence, and camera poses from input images to guide the prediction of 3D Gaussians.
@@
Additionally, several methods build upon DUSt3R [1] for pose-free 3D reconstruction. DUSt3R itself, as a pioneering feed-forward method, utilizes a transformer-based architecture to regress 3D point maps directly from image pairs. Spann3R [62] augments DUSt3R with a spatial memory network, allowing multiview inputs and improving efficiency by eliminating global alignment. However, Spann3R’s sequential processing introduces error accumulation in reconstruction. Fast3R [59] overcomes this limitation by introducing a global fusion transformer, processing multiple views simultaneously and significantly enhancing reconstruction quality. Conversely, CUT3R [3] refines sequential reconstruction by maintaining and incrementally updating a persistent internal state that encodes scene content.
@@
3.4 Dynamic 3D Reconstruction
Compared to static scene reconstruction, dynamic scene reconstruction [198] poses significant challenges mainly due to the presence of moving objects, changing viewpoints, and temporal variations in scene geometry. Extending feedforward 3D reconstruction for dynamic scenarios mainly involves robust pose estimation to mitigate moving object interference, together with dynamic area segmentation for updating changing environments.
@@ 3.4 Dynamic 3D Reconstruction
To jointly resolve pose and obtain a point cloud in canonical space, Robust-CVD [202] and CasualSAM [203] integrate a depth prior with geometric optimization to estimate a smooth camera trajectory, as well as detailed and stable depth and motion map reconstruction. Most recently, MegaSaM [204] further improves pose and depth accuracy by combining the strengths of several prior works, including DROID-SLAM [205], optical flow [206], and a monocular depth estimation model [207], leading to results with previously unachievable quality.
Alternatively, instead of taking advantage of monocular prior models, some methods aim to train a dynamic 3D model from multiview 3D reconstruction models, e.g., DUSt3R [1]. MonST3R [208] estimates pointmap at each timestep and processes them using a temporal sliding window to compute pairwise pointmap for each frame pair with MonST3R and optical flow from an off-the-shelf method. These intermediates then serve as inputs to optimize a global point cloud and per-frame camera poses and intrinsics. Video depth can be directly derived from this unified representation. To speed up the optimization process in MonST3R, DAS3R [209] trains a dense prediction transformer [210] for motion segmentation inference and models the static scene as Gaussian splats with dynamics-aware optimization, allowing for more accurate background reconstruction results. Recent work CUT3R [3] fine-tunes MonST3R [208] on both static and dynamic datasets, achieving feedforward reconstruction but without predicting dynamic object segmentation, thereby entangling the static scene with dynamic objects. Although effective, these methods require costly training on diverse motion patterns to generalize well. In contrast, Easi3R [211] takes an opposite path, exploring a training-free and plug-and-play adaptation that enhances the generalization of DUSt3R variants for dynamic scene reconstruction, achieving accurate dynamic region segmentation, camera pose estimation, and 4D dense point map reconstruction at almost no additional cost on top of DUSt3R. Driv3R [65] further enables dynamic 3D reconstruction in large-scale autonomous driving scenarios by introducing a memory mechanism that supports efficient temporal integration. Besides, it also eliminates the global alignment optimization to reduce computational cost.
@@ 4.2 Evaluation Metrics
For camera pose estimation, RTA (Relative Translation Accuracy), RRA (Relative Rotation Accuracy), and AUC (Area Under Curve) are widely adopted. RTA and RRA measure the relative angular errors in translation and rotation between image pairs, respectively. AUC computes the area under the accuracy curve across different angular thresholds. In point map evaluation, the standard metrics include point cloud Accuracy (or precision), Completeness (or recall), and Chamfer distance. The point cloud accuracy is the average nearest-neighbor distance from each predicted point to the ground-truth surface, indicating how precisely predicted points are placed. The point cloud completeness is the average nearest-neighbor distance from each ground-truth point to the reconstruction, reflecting how fully the ground-truth surface is covered. The Chamfer Distance combines the Accuracy and Completeness scores.
For dynamic point tracking, OA (Occlusion Accuracy), $\sigma^{vis}_{avg}$, and AJ (Average Jaccard) [273] are used together. OA measures the binary accuracy of occlusion predictions; $\sigma^{x}_{avg}$ measures the fraction of points that are accurately tracked within a certain pixel threshold; Average Jaccard considers both occlusion and prediction accuracy.

### v3

@@ 2.5.2 Generative View Synthesis
Regression-based methods work well for view interpolation, while it struggles with view extrapolation, e.g., estimating unseen regions of the scene. In contrast, generative methods instead leverage generative models to synthesize realistic novel views based on learned data distributions, enabling view extrapolation even from a single input image.
@@
Earlier works primarily used transformer-based autoregressive models [124, 149]. For example, GFVS [150] approaches novel view synthesis by sampling target images from a learned distribution conditioned on a source image and camera transformation, where the distribution is modeled autoregressively with a transformer. ViewFormer [151] further extends single-view NVS of GFVS to multiview NVS through a proposed branching attention. Recently, latent diffusion models [152] have been widely used in novel view synthesis due to their ability in generating high-resolution images. These models encode inputs into a latent space using a pretrained variational autoencoder and perform diffusion within the latent space. We categorize them into latent image diffusion models and video diffusion models.
@@
Image Diffusion Models. Zero-1-to-3 [147] leverages the latent diffusion model [152] pretrained for text-to-image generation and replaces text embedding with relative camera poses as conditioning to achieve novel view synthesis as illustrated in Fig. 7(a). ZeroNVS [153] extends Zero-1-to-3 to achieve single-view scene-level novel view synthesis, with an additional field of view as camera parameters to solve the scale ambiguity. However, these methods still face challenges in generating consistent novel views. SyncDreamer [154] addresses this by initializing a diffusion model with Zero-1-to-3 weights and modeling the joint distribution of multiview images. Zero123++[12] arranges six views in a single image for improved multiview representation modeling. ViewDiff [155] also exploits diffusion priors, with an emphasis on multiview consistency in real-world data. Consistent123 [156] combines Zero-1-to-3 and stable diffusion to provide case-aware diffusion priors to ensure multiview-consistency, while ConsistNet [157] performs parallel viewpoint-specific diffusion and aligns the generated images accordingly to enforce multiview geometric consistency. MVDream [88] proposes a multiview diffusion model that leverages both 2D and 3D data, combining the generalizability of 2D diffusion models with the consistency of 3D renderings. With similar multiview diffusion architectures, CAT3D [158] & SEVA [159] and CAT4D [160] substantially advance the performance for static and dynamic view synthesis.
@@
Video Diffusion Models. Video diffusion models [161, 162] have achieved impressive performance in video synthesis and are believed to implicitly capture 3D structures. Building on this capability, recent approaches have explored leveraging video diffusion priors to generate multiview images for high-quality 3D reconstruction. For example, ReconX [148], shown in Fig. 7(b), uses the generative prior of large pretrained video diffusion models [162] to synthesize novel views. It encodes extracted point clouds as 3D structural conditions, ensuring multiview consistency in the generated novel views. Instead of directly using point clouds as condition, ViewCrafter [8] renders point cloud into images as conditions of the video diffusion model [162] to enable consistent and accurate novel view synthesis. MultiDiff [163] leverages a single reference image and a predefined target camera trajectory as conditions of diffusion models, with depth cues to encourage consistent novel view synthesis. More recently, DifFusion3D+ [164], SpatialCrafter [165], and GenFusion [166] have bridged reconstruction and generation by using video diffusion models as scene reconstruction refiners, enabling both artifact removal and scene content expansion.
@@ 3 Tasks & Applications
The preceding methods section highlights the core architectures of feed-forward models, which also include the corresponding basic tasks of novel view synthesis or static 3D reconstruction. In contrast, this section extends beyond the basic tasks and explores the wide variety of applications enabled by feed-forward models.  [^1] We do not cover tasks and applications centered on depth, as this is an independent research line, even though depth is sometimes an output of feed-forward models.
3.1 Camera Pose Estimation
Feed-forward models that predict pointmaps enable efficient recovery of camera parameters without the need for explicit multi-stage SfM pipelines. DUSt3R demonstrates how pose estimation tasks can be naturally derived from pointmap output. Specifically, camera intrinsics such as the focal length can be solved by minimizing reprojection errors of pointmaps. The relative camera motion between two views can be recovered either by first performing pixel matching and intrinsics estimation to compute the essential matrix [167], or more directly by aligning the predicted pointmaps across views through Procrustes alignment [168], which yields rotation and translation up to scale. To improve robustness against noise and outliers, pointmap alignment can also be integrated with PnP-RANSAC [169, 167, 170].
Instead of relying on pointmaps to derive camera parameters, another line of research [171, 3, 2] aims to predict camera parameters directly in a feed-forward manner. Typically, a learnable camera token is prepended to the image tokens, which allows for interaction with other views to capture image-level information, e.g., camera motions. On the other hand, most of them anchor their pose estimation to a reference viewpoint like the first frame, which is an inductive bias that can lead to inferior results when the reference is suboptimal. $\pi^{3}$ [172] instead proposes to estimate the relative camera pose between views, achieving superior performance of camera pose estimation and benefiting other associated tasks such as pointmap estimation.
3.2 Pose-Free 3D Reconstruction & View Synthesis
@@
The development of feed-forward models has enabled the reconstruction of 3D scenes from unposed images or videos without the need for per-scene optimization. FlowCam [173] uses a single-view feed-forward NeRF combined with optical flow to estimate poses and fuse multiview point maps for NeRF-based 3D reconstruction. CoPoNeRF [174] builds and refines 4D correlation maps from image pairs to estimate flow and poses, enabling color and depth rendering.
@@
To extend these approaches to 3DGS, GGRt [93] employs PixelSplat [5] for predicting viewpoint-specific 3D Gaussian maps and introduces a pose estimation module that jointly optimizes camera poses alongside Gaussian predictions. PF3plat [175] proposes a coarse-to-fine strategy, estimating depth, confidence, and camera poses from input images to guide the prediction of 3D Gaussians.
@@
Additionally, several methods build upon DUSt3R [1] for pose-free 3D reconstruction. DUSt3R itself, as a pioneering feed-forward method, utilizes a transformer-based architecture to regress 3D point maps directly from image pairs. Spann3R [62] augments DUSt3R with a spatial memory network, allowing multiview inputs and improving efficiency by eliminating global alignment. However, Spann3R’s sequential processing introduces error accumulation in reconstruction. Fast3R [59] overcomes this limitation by introducing a global fusion transformer, processing multiple views simultaneously, and significantly enhancing reconstruction quality. Conversely, CUT3R [3] refines sequential reconstruction by maintaining and incrementally updating a persistent internal state that encodes scene content.
@@
3.3 Dynamic 3D Reconstruction & View Synthesis
Compared to the basic task of static scene reconstruction, dynamic scene reconstruction [181] poses significant challenges mainly due to the presence of moving objects, changing viewpoints, and temporal variations in scene geometry. Extending feedforward 3D reconstruction for dynamic scenarios mainly involves robust pose estimation to mitigate moving object interference, together with dynamic area segmentation for updating changing environments.
@@ 3.3 Dynamic 3D Reconstruction & View Synthesis
To jointly resolve pose and obtain a point cloud in canonical space, Robust-CVD [185] and CasualSAM [186] integrate a depth prior with geometric optimization to estimate a smooth camera trajectory, as well as a detailed and stable depth and motion map reconstruction. Most recently, MegaSaM [187] further improves pose and depth accuracy by combining the strengths of several prior works, including DROID-SLAM [188], optical flow [189], and a monocular depth estimation model [190], leading to results with previously unachievable quality.
Alternatively, instead of taking advantage of monocular prior models, some methods aim to train a dynamic 3D model from multiview 3D reconstruction models, e.g., DUSt3R [1]. MonST3R [191] estimates the pointmap at each timestep and processes them using a temporal sliding window to compute pairwise pointmap for each frame pair with MonST3R and optical flow from an off-the-shelf method. These intermediates then serve as inputs to optimize a global point cloud and per-frame camera poses and intrinsics. Video depth can be directly derived from this unified representation. To speed up the optimization process in MonST3R, DAS3R [192] trains a dense prediction transformer [193] for motion segmentation inference and models the static scene as Gaussian splats with dynamics-aware optimization, allowing for more accurate background reconstruction results. Recent work CUT3R [3] fine-tunes MonST3R [191] on both static and dynamic datasets, achieving feedforward reconstruction but without predicting dynamic object segmentation, thereby entangling the static scene with dynamic objects. Although effective, these methods require costly training on diverse motion patterns to generalize well. In contrast, Easi3R [194] takes a different path, exploring a training-free and plug-and-play adaptation that enhances the generalization of DUSt3R variants for dynamic scene reconstruction, achieving accurate dynamic region segmentation, camera pose estimation, and 4D dense point map reconstruction at almost no additional cost on top of DUSt3R. Driv3R [65] further enables dynamic 3D reconstruction in large-scale autonomous driving scenarios by introducing a memory mechanism that supports efficient temporal integration. Besides, it also eliminates the global alignment optimization to reduce computational cost.
@@ 3.4 Image Matching
3.5 3D-Aware Image Synthesis
Feed-forward 3D models can be naturally applied to achieve 3D-aware image synthesis by combining 3D representation and generative models (e.g., GAN or diffusion models).
Several earlier methods employ voxel-based representations (e.g., PlatonicGAN [212]) or 3D feature representations (e.g., HoloGAN [213] and BlockGAN [214]). However, these approaches often suffer from limited multiview consistency. To address this, GRAF [215] introduces a generative radiance field as a 3D representation, significantly improving consistency across different viewpoints. PiGAN [216] leverages implicit neural representations with periodic activation functions to model scenes as view-consistent radiance fields. Subsequently, GIRAFFE [217] constructs compositional generative radiance fields for scene representations, enabling controllable image synthesis. StyleNeRF [218] combines NeRF-based 3D representations with a style-based generative model for high-resolution, 3D-consistent image synthesis. EG3D [43] introduces an explicit-implicit triplane representation to achieve efficient and high-quality 3D-aware image synthesis. Due to the high computational cost of volume rendering in implicit NeRF-based scene representations, Hyun et al. propose GSGAN [219], which replaces NeRF with 3D Gaussian Splatting (3DGS), enabling more efficient scene rendering through rasterization-based splatting. To stabilize the training of 3DGS-based 3D-aware image synthesis, GSGAN introduces hierarchical Gaussian representations, enabling coarse-to-fine scene modeling.
With the emergence of diffusion models, a series of works [220, 221, 222, 223] make use of existing 2D diffusion backbones for 3D-aware synthesis by incorporating geometry priors in the form of 3D radiance fields. Instead of predicting NeRF, DiffSplat [224] proposes to generate 3D Gaussian parameters by fine-tuning image diffusion models with structured 3DGS representations. To mitigate the difficulty of 3DGS prediction, Bolt3D [9] employs a diffusion model to predict pointmaps, which are further decoded into 3DGS parameters in a feed-forward manner.
3.6 Camera-Controlled Video Generation
Camera-controlled video generation can be viewed as a natural extension of feed-forward 3D-free view synthesis into video generation. To enable camera pose control in the video generation process, MotionCtrl [225], CameraCtrl [226], I2VControl-Camera [227] inject the camera parameters (extrinsic, Plücker embedding, or point trajectory) into a pretrained video diffusion model. Building upon this, CamCo [228] integrates epipolar constraints into attention layers, while CamTrol [229], NVS-Solver [230], and ViewExtrapolator [231] leverage explicit 3D point cloud renderings to guide the sampling process of the video diffusion models in a training-free manner. AC3D [232] carefully designs the camera representation injection to the pretrained model. ViewCrafter [8], Gen3C [203], and See3D [233] fine-tuned video diffusion models on point cloud renderings to enable better novel view synthesis. VD3D [234] enables camera control to transformer-based video diffusion models. Beyond static scenes, CameraCtrl II [235], and ReCamMaster [236] enable camera-controlled video generation on dynamic scenes by conditioning the video diffusion models on camera extrinsic parameters, while TrajectoryCrafter [237] also enables dynamic scene view synthesis by conditioning the video diffusion models on dynamic point cloud. Several recent works have advanced beyond single-camera scenarios: CVD [238], Vivid-ZOO [239], and SynCamMaster [240] develop frameworks for multi-camera synchronization.
3.7 3D Understanding
@@ 4.2 Evaluation Metrics
For camera pose estimation, RTA (Relative Translation Accuracy), RRA (Relative Rotation Accuracy), and AUC (Area Under Curve) are widely adopted. RTA and RRA measure the relative angular errors in translation and rotation between image pairs, respectively. AUC computes the area under the accuracy curve across different angular thresholds. In point map evaluation, standard metrics include point cloud Accuracy (or precision), Completeness (or recall), and Chamfer distance. The point cloud accuracy is the average nearest-neighbor distance from each predicted point to the ground-truth surface, indicating how precisely predicted points are placed. The point cloud completeness is the average nearest-neighbor distance from each ground-truth point to the reconstruction, reflecting how fully the ground-truth surface is covered. The Chamfer Distance combines the Accuracy and Completeness scores.
For monocular depth estimation, people usually calculate the absolute relative error $|y-\hat{y}|/y$ where $y$ is the ground-truth and $\hat{y}$ is the prediction, and the percentage of inlier points $\delta<T$, which is the percentage of predicted depths within a T-factor of true depth. For multiview depth estimation, several metrics are usually reported including: 1) Accuracy, which measures the smallest Euclidean distance from the prediction to the ground-truth surface; 2) Completeness, which measures the smallest Euclidean distance from the ground truth to prediction; 3) Overall, which is the mean of Accuracy and Completeness, equivalent to the Chamfer distance. Notably, scale ambiguity is an unavoidable issue for monocular depth estimation. A common practice is to perform scale alignment during evaluation, such as median scaling or least-squares fitting.
For dynamic point tracking, OA (Occlusion Accuracy), $\sigma^{vis}_{avg}$, and AJ (Average Jaccard) [286] are used together. OA measures the binary accuracy of the occlusion predictions; $\sigma^{x}_{avg}$ measures the fraction of points that are accurately tracked within a certain pixel threshold; Average Jaccard considers both occlusion and prediction accuracy.